### v1

# A detailed dive into fitting strategies for GRB afterglows with contamination: A case study with kilonovae

Abstract
Observations of gamma-ray burst afterglows have begun to readily reveal contamination from a kilonova or a supernova. This contamination presents significant challenges towards traditional methods of inferring the properties of these phenomena from observations. Given current knowledge of kilonova and afterglow modelling, observations —as expected— with near-infrared bands and at early observing times provide the greatest diagnostic power for both observing the presence of a kilonova and inferences on its properties in gamma-ray burst afterglows. However, contemporaneous observations in radio and X-ray are critical for reducing the afterglow parameter space and for more efficient parameter estimation. We compare different methods for fitting joint kilonova and afterglow observations under different scenarios. We find that ignoring the contribution of one source (even in scenarios where the source is sub-dominant) can lead to significantly biased estimated parameters but could still produce great light curve fits that do not raise suspicion. This bias is also present for analyses that fit data where one source is “subtracted”. In most scenarios, the bias is smaller than the systematic uncertainty inherent to kilonova models but significant for afterglow parameters, particularly in the absence of high-quality radio and X-ray observations. Instead, we show that the most reliable method for inference in any scenario where contamination can not be confidently dismissed is to jointly fit for both an afterglow and kilonova/supernova, and showcase a Bayesian framework to make this joint analysis computationally feasible.
keywords:
methods: data analysis – transients: gamma-ray bursts – transients: neutron star mergers – transients: supernovae1 Introduction
Gamma-ray bursts (GRBs) are broadly classified as either short or long depending on the duration they emit of their energy, T; short for T and long for longer durations. Following the high-energy gamma-ray emission, the relativistic jet collides with the interstellar medium (ISM), producing a broadband afterglow powered (primarily) by synchrotron radiation (Sari et al., 1998). Depending on the type of progenitor system, one may also observe a thermal transient in conjunction with the afterglow, namely a kilonova or supernova. The former are associated with GRBs from compact object merger progenitors and are powered by the radioactive decay of rapid capture (r-process) elements (Lattimer & Schramm, 1974; Li & Paczyński, 1998). Meanwhile the latter are associated with core-collapse of massive stars and are powered (largely) by nickel decay (Arnett, 1982).
The smoking gun multi-messenger gravitational-wave and electromagnetic observations of binary neutron star (BNS) merger, GW170817 and the associated short-duration gamma-ray burst, GRB170817A, confirmed that binary neutron star mergers are the progenitors of some short GRBs (Abbott et al., 2017b; Savchenko et al., 2017; Goldstein et al., 2017). A dedicated follow-up campaign would later reveal the presence of a kilonova, AT2017gfo, confirming BNS mergers as a site of r-process nucleosynthesis (e.g., Abbott et al., 2017a; Villar et al., 2017a, b; Cowperthwaite et al., 2017; Valenti et al., 2017; Tanvir et al., 2017). Although GRB170817A seemingly provides support for classifying GRBs by their duration and progenitor, recent peculiar GRBs like GRB211211A (Rastinejad et al., 2022; Troja et al., 2022) and GRB230307A (Levan et al., 2023; Yang et al., 2024), which are of long duration but appear to be accompanied by a kilonova, alongside previous examples of failed classification based on duration (e.g., Levesque et al., 2010; Ahumada et al., 2021), suggest GRB classification is not straightforward.
Further understanding the nature of the GRB progenitor requires us to accurately determine the properties of the associated kilonova/supernova and afterglow from different observations. This requires developing accurate models for these phenomena and robust inference techniques. For example, parameter estimation of kilonova ejecta properties can play a critical role towards the study of neutron star (NS) post-merger evolution (e.g., Sarin & Lasky, 2021), constraints on the nuclear equation of state (e.g., Raaijmakers et al., 2021), and better understanding of nuclear physics in kilonovae (e.g., Barnes et al., 2016; Metzger, 2017; Hotokezaka & Nakar, 2020).
As the recent observations of GRB211211A and GRB230307A indicate, afterglow and kilonova observations can occur coincidentally, where the observed data is a superposition of the two individual transient emission mechanisms. Disentangling the contaminated signal is crucial to ensure accurate estimation of parameters and correct physical interpretations. Therefore, it is necessary to establish robust methods for inferring properties from observations. Such efforts are complementary to studies such as Ragosta et al. (2024) that investigate observing strategies in the run-up to the Legacy Survey of Space and Time (LSST) (LSST Science Collaboration et al., 2009).
In this paper, we highlight the pitfalls of different strategies for parameter estimation of potentially contaminated afterglow and kilonova signals and showcase a more robust framework for parameter estimation. In particular, we compare the most common analysis technique where the afterglow and/or kilonova is fit separately or on data where the result of fit is “subtracted” (e.g. Jin et al., 2016; Troja et al., 2019; Rastinejad et al., 2022), to the use of a combined model to fit the kilonova and afterglow simultaneously (e.g. Levan et al., 2023). While we focus here on kilonova and gamma-ray burst afterglows, the general techniques described here are directly applicable to other scenarios where two (or more) emission mechanisms appear together.
This paper is structured as follows, in Sec. 2, we describe the emission models for the afterglow and kilonova and our model combining the two emission mechanisms. In Sec. 3, we show lightcurves for a population of afterglows and kilonovae, discuss the regime where observations are most likely to be contaminated, and identify wavelengths and epochs with high diagnostic power. We then go on to compare the methods that can be used to disentangle the transients and infer their properties in Sec. 4, further highlighting the importance of different observing bands and the impact they have on distinguishing between kilonovae and afterglows using photometry alone, and conclude in Sec. 5.
2 Models
Gamma-ray burst afterglows are well described by synchrotron emission produced from the interaction of a relativistic jet with the surrounding interstellar medium. Observations of GRB170817A confirmed the notion that such jets are structured, which could include an inhomogeneous distribution of energy within the jet cone or emission wings that extend beyond the ultra-relativistic core. However, the exact details of jet-structure are uncertain (Granot et al., 2002; Cao et al., 2023). For simplicity, we only consider “tophat” structured afterglows in this work following Lamb et al. (2018), which assumes the jet is a narrow ultra-relativistic core defined by an opening angle , outside of which the energy drops to zero. We note that the choice of a “tophat” structure is likely to well describe a bulk of the parameter space where contamination from a kilonova is most likely to occur: GRBs viewed on axis.
Details of kilonova modelling are far less certain. We consider two commonly used models in the literature; a multi-component kilonova model following Villar et al. (2017a), which we refer to as a ‘two_component_kilonova’ model and the ‘two_layer_stratified_kilonova’ model following Hotokezaka & Nakar (2020). The bulk of the ingredients in these commonly used models are largely similar, such as the input heating rates (Korobkin et al., 2012). However, the models differ most significantly in their treatment of thermalization, with the former following Barnes et al. (2016), and the latter following an independent treatment following Hotokezaka & Nakar (2020). These models also differ in their treatment of the dynamical evolution of the ejecta, with the former following an “arnett” model for supernovae (Arnett, 1982), with appropriate modifications to account for r-process nucleosynthesis (Korobkin et al., 2012; Barnes et al., 2016) and two independent components. The latter model assumes the total mass to be distributed between shells undergoing homologous expansion.
In Fig. 1, we plot the 90% credible intervals for random draws of the models following priors motivated by inferences of AT2017gfo for various observational bands. The models show the largest discrepancies when viewed at higher frequency wavelengths and at later times. The two-component model generally shows greater variation and later decay times than the ’two_layer_stratified’ model. We also note that these models do not capture viewing-angle dependence. However, the effect of viewing angle on kilonova lightcurves is likely significantly smaller than the viewing angle effect on afterglows.
Kilonova lightcurves are far from certain, and inferences of parameters are extremely sensitive to the model choice and the inherent systematics (Sarin & Rosswog, 2024). However, given recent works highlight the important role of the dynamical evolution of the ejecta in shaping the lightcurve (Sarin et al., 2022; Fryer et al., 2023), we focus on the ‘two_layer_stratified_kilonova’ model for the remainder of this work as it considers the dynamics in a more self-consistent manner.
The recent observations of GRB211211A and GRB230307A demonstrate that both kilonovae and afterglows can occur at the same time and create a marked effect on the overall light curve. Meanwhile, numerical simulations have also shown that interaction of the jet with the kilonova ejecta can change the properties of the ejecta, creating significant observable differences to the light curve (Klion et al., 2021; Nativi et al., 2021). For our analysis, we assume that the kilonova and afterglow emission are independent and the observed emission is a linear combination. This simplistic approach is likely unsuitable for on-axis observers, where jet-ejecta interactions likely make the ejecta more energetic and, therefore, ‘bluer’. However, this is not expected to change our conclusions about the ideal methodology for inferring properties, we discuss limitations of our modelling in greater detail in Sec. 5.
3 Contamination regime
The specific time of observations, bandpasses, and model parameters all ultimately impact the afterglow or kilonova lightcurve. We first investigate the parameter space where a kilonova is likely to contaminate observations of gamma-ray burst afterglows. We treat on and off-axis afterglow lightcurves separately in this section due to their distinct appearances. Moreover, the former is more likely to be associated with a prompt gamma-ray burst and accompanied by X-ray and radio observations, alleviating some of the modelling concerns as we discuss later.
As the distinct behaviour of on-axis and off-axis observers suggests, the degree of contamination from the kilonova to the afterglow is ultimately sensitive to the afterglow peak timescale. This timescale, in turn is sensitive to the core opening angle , jet energy , and interstellar medium density (Granot et al., 2002). Similarly, the kilonova peak timescale (dependent on the opacity, mass, and velocity) also plays a role. However, kilonova lightcurves have more significant variation between observational bandpasses, such that the dependence on the peak timescale itself is minimal compared to the afterglow.
To demonstrate the possible degree of contamination, we randomly select 5000 kilonova and afterglow parameter sets from priors motivated by inferences of the kilonova, AT2017gfo (Villar et al., 2017a), and afterglow of GRB170817A (Cao et al., 2023) all at the same distance of 44.6 Mpc. We do this twice for the afterglow, once to represent an on-axis observation scenario and once to represent an off-axis scenario. The 90% credible intervals for resulting lightcurves are shown in Fig. 2, and we specify the parameters and priors of the afterglow and kilonova models in Tab 1 and 2, respectively. In Fig. 3, we explicitly show how the fractional contribution of flux from the kilonova, relative to the combined flux as expected from Fig. 2, evolves with time. We include all LSST filters and plot the median and credible interval for the kilonova contribution fraction, highlighting the degree of contamination from the afterglow at different wavelengths. From the figures, we immediately see strong dependence on the viewing angle , with on-axis systems likely suffering from significant contamination at all wavelengths. Meanwhile, systems viewed off-axis are only susceptible to contamination at late-times (). As a specific example, we can consider the case of GW170817, which was observed at least off-axis, and there the kilonova had no discernible impact on the afterglow.
| Parameter | Distribution | Range |
| Redshift, | Fixed | 0.01 |
| Viewing angle, | Sine | (0, /2) rad |
| Jet half opening angle, | Uniform | (0, 0.1) rad |
| Isotropic jet energy, | Uniform | (44, 54) |
| ISM density, | Uniform | (-5, 2) |
| Accelerated electron fraction, | Fixed | 1 |
| Electron spectral index, | Fixed | 2.3 |
| Fraction of energy in electrons, | Fixed | -1.25 |
| Fraction of energy in magnetic field, | Fixed | -2.5 |
| Initial Lorentz factor, | Fixed | 1000 |
| Parameters | Distribution | Range |
| Redshift, | Uniform | (0.001, 0.1) |
| Ejecta mass, | Uniform | (0.01, 0.05) |
| Inner shell velocity, | Uniform | (0.05, 0.2) |
| Outer shell velocity, | Uniform | (0.3, 0.5) |
| Grey opactity, | Uniform | (1, 30) cm2/g |
| Density profile power law index, | Uniform | (3.1, 8) |
Investigating the specific on-axis afterglow case further. We generally find that the possibility of significant kilonova contamination strongly depends on the overall brightness of the afterglow. In particular, for our choice of fiducial distance, we find significant contamination, provided the afterglow is brighter than at post GRB, in the lsst-r band. Inspecting Figs. 2a and 3a, we see that bandpasses at higher effective wavelengths are more likely to have significant contamination, a direct consequence of the competition between the non-thermal GRB and thermal (but rapidly cooling) kilonova spectrum. In particular, almost all near-infrared (NIR) to optical wavelengths have significant periods where the kilonova flux density contributes almost entirely to the total flux density across all our random draws. However, at bluer wavelengths e.g., UVOT:uvw1 filter, the kilonova contribution does not exceed 0.5% at any time or for any model, suggesting that this filter offers the best opportunity (at near optical wavelengths) to probe just the afterglow. The 90% credible intervals in Fig. 2a show that contamination is likely in all wavelengths, however at lower frequency bandpasses where the kilonova evolves the slowest, it is possible to identify the transient as a bump superimposed on the afterglow lightcurve. We conclude that wavelengths from NIR to lsst-r hold the greatest diagnostic power for kilonova detection and subsequent parameter estimation, with the optimal observation window ranging between approximately . However, we do emphasize that these statements are based on the broad interpretation of our randomly sampled population, and specific cases may be significantly different.
Much more so than the on-axis scenario, contamination between an off-axis afterglow and a kilonova is strongly sensitive to the afterglow peak timescale and brightness, with earlier and brighter peaks more likely to contribute to signal overlap. For our choice of distance, we find significant contamination to only occur where the peak flux density is greater than and occurs at a time before post GRB in the lsst-r band. The impact of observational bandpasses on kilonova contamination for an off-axis GRB is shown in Figs. 2b and 3b. We draw attention to the certainty of kilonova domination at early times. In the NIR bands, we see that the kilonova contributes almost all of the observable flux up to at least post GRB, but this window can decrease to as little as in the UV. After such times we see the kilonova contribution fall rapidly due to both the onset of kilonova decay and the afterglow rising to peak, as depicted in Fig. 2b. This further emphasises the high diagnostic power of NIR bands and the importance of early observation times for kilonova detection and parameter estimation. Although this bodes well for the kilonova, it may hinder the afterglow recovery, especially as in an off-axis scenario we are less likely to have X-ray or radio coverage, which is more likely triggered by the detection of a GRB. As we show in more detail in Sec. 4.3, the lack of the radio and X-ray data weakens our ability to accurately infer the parameters of the kilonova and afterglow.
4 Simulations
We now explore different strategies for fitting observations under different contamination/observational scenarios. In particular, we apply different fitting techniques to simulated flux density data sets, ensuring we know all aspects of the underlying model, noise, and true input parameters. The data represents contamination scenarios involving both on- and off-axis afterglows and situations with and without X-ray and radio observations. The afterglow and kilonova are fit individually, using the same transient model as the input, and jointly, using a combined transient model. Individual kilonova fits are also made to data that we first independently subtract the inferred afterglow from. By comparing the estimated values and fits of these different scenarios with the input values used to simulate the relevant data set, we aim to understand what the ideal method and observational practices are for robust parameter estimation and interpretation.
The dominant cost of fitting for all our scenarios was the cost of the afterglow model, where each evaluation of the model could take up to . To reduce this cost, we first develop a machine-learning surrogate model for the afterglow model using a multi-layer perceptron (MLP) neural network, with the scikit-learn package. The model hyperparameters are chosen such that the network is composed of three hidden layers, each containing 300 perceptrons, with the rectified linear unit function as the non-linear activation function and adam as the optimizer (Kingma & Ba, 2014). We use 743,985 samples generated with our actual afterglow model, with a train-test split of 0.15/0.85. Samples are drawn from the unconstrained prior distributions of the tophat afterglow with only redshift, and electron participation fraction, as fixed parameters. To ensure a representative training set, we distinguish between viewer alignments (on-axis and off-axis afterglows) and split the frequency range into nine logarithmic bins. Although redshift is fixed in our samples, we account for distance variations by applying a scaling correction to the flux density within redback for other specified values. We pre-process training samples in a similar manner to Sarin et al. (2021); Boersma & van Leeuwen (2023). This is done by first taking the base 10 logarithm of every parameter, except , and applying the standard scaler function from scikit-learn (Pedregosa et al., 2011). Similarly, we pre-process the target flux density values by taking their natural logarithm and applying the standard scaler function. After training, we find the test set returns a score of 0.99, indicating the emulator is suitably accurate relative to the original ’tophat_redback’ model. We have further validated the surrogate model by performing full simulations where we generate data with the true numerical model and perform inference with the surrogate. Timing the generation of flux density values on consistent arrays, we find the emulator is faster than the original model by up to 1-2 orders of magnitude depending on the size of the output frequency and time arrays. We make this model publicly available through the Redback package.
We explore two primary synthetic data, both representing scenarios where there is significant contamination between a kilonova and an afterglow and an additional scenario where the kilonova is sub-dominant (Sec. 4.4). The first scenario in Sec. 4.1 reflects the observations of a system viewed on-axis, and the second in Sec. 4.2 reflects those of a system viewed off-axis. In each case, we generate 100 synthetic data points with normally distributed noise, at a redshift of , and an extinction of . The observational data is randomly generated for the bands presented in the previous sections as well as at frequencies of and , to represent radio and X-ray emission. Between the on-axis and off-axis significant contamination scenarios, we only change three parameters: , , for the on-axis system and , , for the off-axis system. The remainder of the afterglow injection parameters are the same between the two scenarios with values of , , , , , . We also keep the kilonova the same with injection values of , , , and .
For the two different sets of data, we compare different strategies to fit the observations and infer the properties of the kilonova and afterglow. 1) We explore fitting the afterglow and kilonova models to the data independently, fitting only one transient at a time to the contaminated data. 2) Joint fitting, where we simultaneously fit the data with the combined model and 3) A strategy commonly used, where we fit the kilonova to the subtraction of the best-fitting afterglow model from an independent afterglow-only fit. We fit the observations using the same models as the input with the strategies outlined above. The fits are performed through Redback (Sarin et al., 2024) using a combination of the Dynesty (Speagle & Barbary, 2018) and emcee (Foreman-Mackey et al., 2013) samplers wrapped within Bilby (Ashton et al., 2019).
Our fits for the different scenarios are presented in Fig. 4, where each subplot shows the predicted lightcurves generated from the posterior distributions as solid lines, with colours corresponding to the observation band. For comparison, we overplot the expected curves based on the injection values of the afterglow and kilonova components as dashed and dotted lines, respectively. Figs. 5 and 6, show the posterior distributions of relevant parameters under different fitting strategies.
In Sec. 4.3, we investigate how the availability of observational data impacts the ability to disentangle the transients. Since the kilonova does not emit at X-ray and radio frequencies, observations of the afterglow in these bands are particularly interesting as they can constrain the afterglow parameter space during inference to make the overall fitting process more efficient. To highlight their importance, individual afterglow and joint fits are made to the same two data sets as in the previous sections, but with the X-ray and radio observations removed.
4.1 Scenario 1 - On-axis afterglow with significant kilonova contamination
As discussed in Sec. 3, an on-axis GRB system is likely to suffer from significant kilonova contamination. As the cases of GRB211211A and GRB230307A already prove, they are also the most observationally relevant. We first focus on the strategy where the kilonova is fit independently to the data. The fitted lightcurve from this strategy is shown in Fig. 4 (panel B), which highlights the poor recovery relative to the true input. However, we emphasize that in real observation scenarios, true parameter values are not known, so one could not directly assess the accuracy of inference in this manner. If we disregard the true input curves, the posteriors in Fig. 4 appear to reasonably match the relevant data points for the transient being fit, suggesting a ‘good’ fit to the data that may lead to the conclusion that the inferred parameters are correct.
In Fig. 5, we show violin plots for relevant kilonova parameters from a fit to the data with just a kilonova model in orange, with injection values marked by . Despite producing good fits by eye to the data, our inferred parameters are significantly biased. In particular, and the ejecta mass where the bias is and , respectively (relative to the statistical errors), with no posterior overlap with the true input. Given that these parameters are the most important for interpreting a kilonova, the error seen here suggests that solely fitting the kilonova is not a good strategy.
As shown in Fig. 4 (panel A), the independent fit to the afterglow is a more viable strategy in producing a better match to the true input. This is largely a product of the fit, including the X-ray and radio data, which provide some “calibration” for the afterglow model as the kilonova model has flux at these wavelengths. This motivates the use of the commonly-used strategy of attempting to isolate the kilonova signal by first subtracting the predicted afterglow, and then performing a fit. We approach this “subtraction” strategy by simply masking the X-ray and radio data points and then subtracting the afterglow contribution with the best-fit model parameters for all other wavelengths. To account for the uncertainty in the subtraction, we associate with each subtracted data point the same error as the original data point.
In Fig. 5, we show the posteriors obtained by fitting the kilonova model to the data, subtracting the best-fit afterglow model (in orange). While these posteriors are somewhat more consistent with the input than the independent, kilonova-only fit (cyan), they are still significantly biased. Notably, again, neither the kilonova ejecta mass nor shell velocity is consistent with the input, highlighting the pitfalls of this commonly used strategy. We note that increasing the size of the (largely arbitrary) uncertainty on the error bars when fitting the afterglow-subtracted data could make the kilonova parameters more consistent with the input. However, from a Bayesian perspective, the “subtraction” approach is incorrect and an example of a misspecified model. In particular, when we first fit the afterglow model to data, to maximise the likelihood, the afterglow will depart from the ‘truth’, especially when radio and X-ray data errors are high or the quality otherwise due to cadence or spectral coverage is low relative to the optical. This biased afterglow recovery is then built into the kilonova parameter estimation. Given the intrinsic kilonova and afterglow models are themselves uncertain (unlike our simulation where we know the input), the subtraction strategy can cause significant issues with interpretations. By contrast, the proper (jointly inferring) Bayesian approach to this problem produces kilonova and afterglow posteriors that are marginalising over the other (afterglow) model as opposed to being conditioned on it, and so they will naturally be consistent with the input (given adequate sampling). Moreover, by marginalising over the other model parameters, it is far more likely to yield robust estimated kilonova parameters, as opposed to the subtraction approach.
To verify that the joint approach works, we show fits to the data in Fig. 4 (panel C), which shows that the independent kilonova and afterglow models are consistent with the input. A full corner plot for the joint recovery can be found in Fig. 10 of the appendix, which confirms an accurate recovery with all true kilonova and afterglow parameters recovered correctly to within 3 of the predicted values. We also emphasize that the joint-fitting approach is not significantly more computationally intensive than independently performing an afterglow-only and, subsequently, kilonova-only fit. And importantly, it does not require extra data processing to “generate” afterglow-subtracted kilonova data.
4.2 Scenario 2 - Off-axis afterglow with significant kilonova contamination
Following the same analysis approaches as Sec. 4.1, we again compare individual, including subtracted, and joint fits for an off-axis GRB system. In Sec. 3 we noted that at early observation epochs, the off-axis afterglow flux density is typically low, which allows the kilonova to dominate the signal in all wavelengths. This suggests that strategies which involve individually fitting the kilonova to the data may be more suited towards the off-axis contamination scenario than the on-axis scenario. The result of our individually fit kilonova component relative to the true inputs, shown in Fig. 4 (panel E), seems to reflect this.
In the lower panel of Fig. 6, we show a violin plot in blue for the individual kilonova recovery. The biases on ejecta mass and shell velocity are reduced in comparison to the on-axis GRB case, discussed in Sec. 4.1. However, the posterior of remains inconsistent with our input. We additionally see that opacity is incorrectly inferred, which is likely due to sensitivity to afterglow contamination around the knee of the lightcurve, which is typically the source of constraints on opacity (alongside the peak time) for the kilonova models used in this work. The inaccuracy of the recovered posteriors despite a seemingly good fit to relevant data points, highlights the need for caution in interpreting fits to subtracted data, and general caution for detailed parameter estimates of quantities with this fitting strategy.
We now look at the individual afterglow fit to data and note that while the kilonova is largely dominant in the lower frequency bandpasses, up to 10 days in lsst-r, this does not necessarily detriment the afterglow recovery. This is again due to the inclusion of uncontaminated radio and X-ray observations which provide a level of calibration to the fit. In the upper panel of Fig. 6, we show violin plots for the key afterglow parameters which we identified as significantly influential to contamination in Sec. 3. Although these properties are recovered at a 3 level (relative to statistical errors), the full recovery shows that there are greater biases on other parameters such as extinction and initial lorentz factor, which are largely informed by the optical/NIR light curve. These inconsistencies again, prompt a review of methods to obtain more accurate predictions so that one may better interpret the afterglow.
If we choose to subtract the best fit afterglow from the data to isolate the kilonova, the resulting inference still fails to correctly recover the input value of ejecta mass. Instead, we may again turn to the use of a combined model. We show the joint fit in panel F of Fig. 4, and note that it provides a good visual fit to the true curves of both afterglow and kilonova components. The full simultaneous inference results show that every parameter is not only recovered correctly, but the predictions have also improved in accuracy. Relative to statistical errors, all afterglow properties are inferred correctly to within and all kilonova properties are within . We include the complete corner plot for all inferred parameters in Fig. 11 in the appendix. These results provide additional evidence that jointly fitting a contaminated signal is the most accurate and reliable method to use for disentangling and interpreting these transients.
4.3 Scenario 3 - Data without X-ray and radio bands
We have surmised that the individual afterglow recoveries thus far benefit from the presence of X-ray and radio observations that remain free of kilonova contamination. Here, we explicitly test this by starting with the same data sets used in Sec. 4.1 and Sec. 4.2. We first remove the X-ray and radio data points, and then reattempt to recover the afterglow component with our different fitting strategies. The result of this individual afterglow fit for the on-axis system, without the extra bandpasses for calibration, predicts most parameter values to be correct to within 3. However, for many parameters the magnitude of the statistical uncertainties and therefore the width of the posterior distributions have increased. This suggests that without X-ray and radio observations to ’guide’ the inference, the precision of property predictions is reduced.
The power of the uncontaminated X-ray and radio bandpasses is further emphasised when we consider the individual afterglow fit for the off-axis system as presented in Fig. 7a. This afterglow fit is much worse than that of the on-axis system, and recovers only , and correctly to within 3. When we observe the predicted lightcurve from the posteriors in Fig. 7a and compare it to the expected lightcurve based on the injection values, the obvious mismatch makes it clear why the majority of inferred properties are incorrect. As the true afterglow lightcurve shape is difficult to determine from the data, and the X-ray and radio observations are not available to provide such information, the initially kilonova dominated data points appear to be fit as an extension of the afterglow signal itself. We saw in Fig. 4 (panel E) from Sec. 4.2 that the equivalent kilonova only recovery also fits these same data points and correctly infers the transient parameters. Since the individual fits contradict each other, if confronted with real data lacking the radio and X-ray observations, it could be difficult to determine which of the individual transient recoveries are correct. This means in the case of an off-axis system, which may have contamination and is more likely missing crucial X-ray and radio observations, individual fits or those performed after subtraction are not reliable.
Given the stark failure of individual fitting methods to the off-axis system, we apply the joint fitting strategy to both the on and off-axis scenarios, and compare our results. In both cases, we see improvement and find that all inferred parameters of the kilonovae and afterglows are consistent with the input values. The use of a combined model allows a correct separation of the data points into their respective transient components, which resolves the problem observed in the individual fit of Fig. 7a. However, we note that while accuracy has improved, the posterior distributions have broadened (as expected) compared to the joint fits from Secs. 4.1 and 4.2, which included X-ray and radio observations. An example of this can be seen in the appendix, where we show the full corner plot from the joint inference of the off-axis GRB system.
Another benefit of obtaining X-ray and radio observations is that it reduces the computational expense of Bayesian inference. Without the uncontaminated afterglow data to provide a ‘guide‘ when fitting, the sampling process is significantly longer as the efficiency of finding suitable draws from the prior is lower than it would be with the X-ray and radio data included. With MCMC methods where the number of posterior draws can be chosen, we found that often longer chains were required to produce satisfactory convergence. In this section, we therefore conclude that obtaining X-ray and radio data holds great value in terms of obtaining higher precision inference and reducing computational expenses. However, where it is not possible to obtain these observations, particularly in off-axis scenarios where prompt emission accompanied by X-ray and radio data is less likely to be detected, we emphasise that a joint fit is most likely the only way to correctly interpret the properties of the kilonova and afterglow. Additionally, we find that joint fits are less susceptible to problems associated with multi-modal distributions when using MCMC sampling. Bayesian inference with the combined model therefore proves to be the most suitable choice across all scenarios.
4.4 Scenario 4 - On-axis afterglow with little kilonova contamination
Thus far, we have shown that significant contamination between an afterglow and a kilonova can cause notable biases in parameter inference if the most common individual fitting strategies continue to be used. Here, we consider the scenario where contamination is much less significant. In this scenario, data is simulated for an on-axis GRB system with our key afterglow parameters changed to , , and , creating a particularly bright transient. The remainder of the injection values for the afterglow and kilonova remain the same as in the previous scenarios. This results in an afterglow dominated situation with very little contamination from the kilonova, which is observable as a small bump on the lightcurve at frequencies less than that of the lsst-z band.
Beginning, as usual, with the individual fits, we show in Fig. 8 the violin plots for the inference of key afterglow parameters (upper pannel in red) and kilonova parameters (lower panel in blue). These plots show that there is heavy bias in both the afterglow and kilonova individual fits, despite the afterglow being the dominant component and containing X-ray and radio data for calibration. By employing Bayesian inference with the joint model, we see that all properties of both transients are consistent with the injection values. The best fit posterior lightcurve for the joint fit is shown in Fig. 9 which makes it easy to see how subtle the contamination is compared to panel C in Fig. 4. Our results emphasise that where contamination is suspected, even if not necessarily obvious, a combined model is required to allow an accurate interpretation of the transients.
5 Conclusions
Detailed observations are beginning to readily reveal contamination of a GRB afterglow with additional components. In this work, we have explored several strategies for fitting such observations and inferring the properties of the individual afterglow and additional component. In particular, we have focused on the scenario of a kilonova and an afterglow motivated by recent observations of long gamma-ray bursts afterglows with infrared excesses consistent with a kilonova (Rastinejad et al., 2022; Levan et al., 2023; Yang et al., 2024).
We first determine the region of parameter space (time, wavelength and afterglow and kilonova properties) where contamination is more likely. In general, kilonova emission is brighter and lasts longer in the lower frequency NIR bands, up to post GRB, in comparison with higher frequency UV bands where the onset of decay can begin as soon as post GRB. In an on-axis contamination scenario, the kilonova will be observed as a bump in the lightcurve, the significance of which is determined by the flux density of the afterglow component. The optimal observation window in this case is approximately . In an off-axis scenario, the kilonova will always initially dominate the signal, which is favourable for kilonova parameter estimation. The takeaway, regardless of viewing angle, is to observe as early as possible and make use of NIR bands if aiming to detect a kilonova, which is consistent with current observational strategy.
We then explored different strategies to fit the data and estimate the parameters of the kilonova and afterglow using synthetic, simulated data where we knew the true input. In many cases where the afterglow and kilonova are fit individually, a predicted lightcurve may look reasonable to the eye but the estimated parameters are completely inconsistent with the true input, highlighting the dangers of this strategy for correct interpretations. By investigating different contamination scenarios, we find that for on-axis observations, the kilonova cannot be recovered correctly using individual model fitting. The more commonly used strategy of subtracting a best-fit afterglow and then fitting a kilonova fares better, but again the kilonova parameters can be biased. We note that this also does not produce a fully marginalised posterior, but instead a posterior on the kilonova parameters conditioned on the afterglow model and best-fit parameters, which is not ideal. We find that the best fitting strategy is to jointly fit both the afterglow and kilonova and have showcased a Bayesian framework (publicly available via Redback) to perform such analyses.
When the afterglow is missing observations in the radio and X-ray bands, it can be difficult to distinguish between data points that are strictly from afterglow contributions and those that are strictly from kilonova contributions. This is especially true for off-axis systems where it is possible for data points from the initially kilonova dominated signal to be fit as if they were an extension of the afterglow. Individual parameter inference is unreliable in these scenarios and is difficult to improve via a subtraction technique due to contradicting fits between the afterglow and kilonova. The only solution here is to again use a joint-fitting approach. However, while X-ray and radio data are not necessary to ensure a successful joint fit, these observations hold the power to decrease computational time by improving sampling efficiency and significantly improve the precision of predictions and, therefore, should remain a priority when observing.
In this paper, we have demonstrated that where there is possible contamination between a kilonova and an afterglow, the use of a joint model for simultaneous transient inference is the ideal methodology for accurate and robust parameter estimation. Joint fitting also proves to be the most reliable method for inference, succeeding in situations where individual fits or subtraction techniques fail. While the combined model has been successful on our synthetic data sets, we recognise that there is significant room for development to make the combined model more realistic. In particular, our combined model implicitly assumes that the afterglow and kilonova emission is uncoupled, contrary to numerical simulations that show that kilonova ejecta properties change due to the traversal of a jet (Klion et al., 2021; Nativi et al., 2021). Furthermore, our analytic models, are limited by the assumption of a constant, grey opacity, which implicitly builds a shape to the overall kilonova lightcurve that numerical simulations of kilonova lightcurve do not always obey. An alternative approach to the semi-analytical models we use would be to use machine-learning based surrogates to model the numerical simulations of kilonovae (e.g., Lukošiute et al., 2022). However, these simulations are performed under the assumptions of local thermal equilibrium (LTE), which is invalid for most kilonova parameters past (Pognan et al., 2023), and can have a significant impact on the lightcurves past this epoch. Few studies have performed non-LTE kilonova modelling due to dramatically higher computational cost and nuclear data required to compute such lightcurves. As this work uses only simulated data, non-LTE effects are not considered, and the simplified semi-analytical formulae for kilonovae are considered the ‘true’ input for our analysis, even past , which could affect our conclusions. Further detailed modelling may also account for an afterglow jet with angular structure. Both such changes may help with the accuracy of parameter estimation, but it is unlikely that they will significantly change our conclusions based on the choice of methodology.
An important caveat to our exploration of the idealised approach to fitting is the consideration that the models themselves are also inherently uncertain. And so while the choice of fitting strategy, e.g. subtraction may bias our results, the bias is (at least in most of our simulated scenarios) smaller than the systematic uncertainty in the kilonova models themselves due to uncertain underlying physics (Sarin & Rosswog, 2024; Brethauer et al., 2024). Therefore, we expect broad (order of magnitude estimates) of kilonova properties from subtraction-based fitting strategies (e.g., Rastinejad et al., 2022; Rastinejad et al., 2024) are likely robust. However, the bias is likely more of a concern for the relatively more certain afterglow parameters. We leave a detailed exploration with real observations for future work. We note that regardless of the size of bias a fit based on individual component fitting or subtraction creates, it is not the ideal quantity one desires when estimating parameters through Bayesian inference; a marginalised posterior distribution, dependent only on the observed data. Instead these alternative approaches implicitly condition the posterior on the choice and parameters of a best-fit afterglow model or some other underlying assumption. This is not ideal when what we truly desire is a posterior marginalising over every other degree of freedom.
Acknowledgements
W. Wallace would like to thank Nordita for hosting her visit for the duration of this project, and acknowledges support for the mobility from a Turing Scheme grant funded by the UK government and awarded through the University of Bath. Through the CRESST II CoOp W.Wallace is also supported by the National Aeronautics and Space Administration (NASA) under award number 80GSFC21M0002. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NASA. N.S acknowledges support from the Knut and Alice Wallenberg foundation through the “Gravity Meets Light” project, and the support through the Nordita Fellowship. Nordita is supported in part by NordForsk.
Data Availability
Analysis and fitting is carried out in python using the open source software package redback (Sarin et al., 2024), with the data generated via the simulation interface of this package. The models used/developed for this project are implemented in redback and the neural network is implemented in redback_surrogates. Both software packages are available on pypi, with documentation at https://redback.readthedocs.io/en/latest/ and https://redback-surrogates.readthedocs.io/en/latest/ respectively.
Appendix
Here we include the full posteriors for different analyses performed in this work.